OpenStudy (anonymous):

Let f(x)= ((3x+1)/(x=2), with x<3 and x cant equal -2

OpenStudy (anonymous):

hmm... i dont seem to understand your equation. are you sure ur trying to divide by x=2?

OpenStudy (anonymous):

Let \[f(x)=\left\{ (3x+1)/(x+2), x<3, x \neq-2 \right\}\] and \[\left\{ (\sqrt{x+1 } , x \ge3 \right\}\]

OpenStudy (anonymous):

Find the points of discontinuity of the function f. Identify each type of continuity.